Journalist Report – March 5th

Sol 12 – Bad weather, good times

Author: Nicolas Wattelle

The weather is definitely not kind to us, it pushes us to take some forced rest. Outside, the wind strongly blows, the rain and hail fall, and snow is to come during the night…

On one hand, technically speaking, it means taking the LOAC (aerosol meter) back in the Hab, because it might not survive under the snow, do maintenance on our HF antenna which is put to trouble by the wind, and hope the batteries are empty until tomorrow morning…

On the other hand, today’s experiments needed an outside intervention (geology, Augmented Reality, HF propagation…), more than just doing maintenance around the Hab, so we were stuck on these points.

So, we decided yesterday’s evening to “advance” our rest day, because of this, also because we did not want to lose efficiency in our upcoming work, we try thus to avoid tiredness: after the maintenance EVA, our afternoon would be declared empty, a time to breath, a time to relax.

We are finishing our second week on Mars, and even if some of the last days were not that intense, we lately feel more frequently tired. Maybe the weight of the simulation, maybe the repetitiveness of actions during the days, maybe the accumulation of little frustrations that are taken individually means almost nothing but when added it gets revealed to day light… Maybe a mix of all of that.

Even if we do not know exactly why, one of our strengths is to know how to overcome it: relax! And guess what, we are particularly good at it!

Our afternoon was thus focused on making us having a great time, cooking, chatting, living good! Some of us played games, others drew and painted, listened to music.

This looked like last Sunday, even if outside the bad weather strikes, inside the Hab it is bright and sunny!

Operations Report – March 5th

Crew 263 Operations Report 05-03-2022

SOL: 12

Name of person filing report: Mathéo Fouchet

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 169.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 85.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes and comments: Nothing to report

Summary of Hab operations:

WATER USE: 15 gallons

Water (static tank): 325 gallons

Water (loft tank): 40 gallons

Water Meter: 01562733 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: Nothing to report

Summary of suits and radios: Today during our morning EVA, one of the EVA participant could not hear his teammates through his earpiece. We do not really understand why this kind of problem is happening with our earpieces because we treat them with great care.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations:

Dual split: Off

SOC: 65% at 6:00pm

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: Would it be possible to have air filters so that I can change the air filter in the lower deck in the Hab?

EVA Report – March 5th

Crew 263 EVA Report 05-03-2022

EVA # 16

Author: Valentine Bourgeois

Purpose of EVA: Atmospheric Experiment Battery Change and reorientation of the antennas.

The atmospheric experiment includes 5 devices: LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a Weather station. The aim of this experiment is to do testing of Martian-Atmospheric Study Equipment and their maintenance by human operators. Performing their maintenance and set-up while wearing spacesuits and following EVA protocols is the purpose of the experiment. All devices are located at the same place.

Batteries for the LOAC, Mega-Ares and Purple Air need to be changed daily. Hence, during this maintenance EVA, the 3 batteries needed to be replaced. As there is a chance of snowing tonight and tomorrow, we also decided to bring the LOAC back to the station as it might be damaged by the snow.

Start time: 10:36

End time: 11:28

Narrative: EVA Participants entered the airlock at 10:36 Depressurization ended at 10:41. EVA participants checked the rover batteries just in case. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 85.2 Battery Charge – 100%

Curiosity: Usage Duration – 169.2 Battery Charge – 100%

The EVA participants arrived at the Atmospheric Experiment location at 10:46. The 2 batteries were replaced. Marine turned off the LOAC and disconnected the battery. Nicolas changed the battery of Mega-Ares and protected it with a second garbage bag. Valentine found some rocks to elevate the battery in case of snow tomorrow. Léa changed the Purple Air battery. The status of the devices was checked.

At 10:52, the participants left the atmospheric experiment location. With two old batteries and the LOAC. They have been placed in front of the air-lock to continue the EVA. We went to the antenna to fix it.

At 10:55, the participants arrived at the antenna location. Because of the wind the antennas have changed their orientation. We fixed it and place some more ropes and rocks to keep them from doing it again.

When we were almost done, Valentine’s earphone broke. She couldn’t hear anything but her teammates where able to hear her. We finished to fix the last rope and went back to the Hab. We took the old batteries and the LOAC with us to go back to the air-lock.

11:23 pressurization over.

Destination: Atmospheric experience site (at intersection between Cow Dung Road and Entrance Road) and the antenna site

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants :Valentine Bourgeois (EVA Leader), Léa Rouverand, Nicolas Wattelle and Marine Prunier

Road(s) and routes per MDRS Map: Entrance Road

Mode of travel: Foot

Sol Summary – March 5th

Crew 263 Sol Summary 05-03-2022

Sol: 12

Summary Title: Bonus rest day

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: As we decided yesterday in the evening, we skipped the morning sport session and slept in. The Morning EVA was planned quite late so we had time to chill and eat pancakes for breakfast together. The EVA was necessary to change the atmospheric experiment batteries, remove one of the devices that do not handle snow well, and re-oriente the HF antenna. We canceled the afternoon EVA because it was very windy and sometimes rainy. It was a very nice rest day, everyone seems to have a regain of energy.

Anomalies in work: None.

Weather: Unsettled.

Crew Physical Status: Very good.

EVA: EVA 16 in the morning, nominal. EVA 17 was cancelled.

Reports to be filed: Sol Summary, EVA Request (n°17), EVA Report (n°16), Operations Report, Journalist Report, HSO Report.

Support Requested: No support requested.

Copyright © The Mars Society. All rights reserved. | Main Site